\begin{thebibliography}{99}
\bibitem{Abbas} Abbas, M. \& Nazir, T. 2013. A new faster iteration process
applied to constrained minimization and feasibility problems. Matematicki
Vesnik, Article.

\bibitem{S} Agarwal, R., O Regan, D. \& Sahu, D. 2007. Iterative
construction of fixed points of nearly asymptotically nonexpansive mappings.
Journal of Nonlinear and Convex Analysis 8: 61-79.

\bibitem{Babu} Babu, G. V. R. \& Prasad, K. N. V. V. V. 2006. Mann iteration
converges faster than Ishikawa iteration for the class of Zamfirescu
operators. Fixed Point Theory and Applications 2006(1).

\bibitem{Vasile} Berinde, V. 2007. Iterative Approximation of Fixed Points.
Springer, Berlin.

\bibitem{Berinde} Berinde, V. 2004. Picard iteration converges faster than
the Mann iteration in the class of quasi-contractive operators. Fixed Point
Theory Appl. 2: 97-105.

\bibitem{Berinde1} Berinde, V. 2004. Comparing Krasnoselskij and Mann
iterations for Lipschitzian generalized pseudocontractive operators.
Proceed. of Int. Conf. On Fixed Point Theory, Univ. of Valencia. Yokohama
Publishers.

\bibitem{Berinde2} Berinde, V. \& Berinde, M. 2005. The fastest
Krasnoselskij iteration for approximating fixed points of strictly
pseudo-contractive mappings. Carpathian J. Math 21(1-2): 13-20.

\bibitem{SS} Chang, S. S., Cho, Y. J. \& Kim, J. K. 2003. The equivalence
between the convergence of modified Picard, modified Mann, and modified
Ishikawa iterations. Mathematical and Computer Modelling 37(9): 985-991.

\bibitem{CR} Chugh, R., Kumar, V. \& Kumar, S. 2012. Strong Convergence of a
new three step iterative scheme in Banach spaces. American Journal of
Computational Mathematics 2: 345-357.

\bibitem{Ciric} Ciric, L. B., Lee, B. S., \& Rafiq, A. 2010. Faster Noor
iterations. Indian Journal of Mathematics, 52(3), 429-436.

\bibitem{GK} G\"{u}rsoy, F. \& Karakaya, V. 2014. Some Convergence and
Stability Results for Two New Kirk Type Hybrid Fixed Point Iterative
Algorithms. Journal of Function Spaces doi:10.1155/2014/684191.

\bibitem{KMSP} G\"{u}rsoy, F., Karakaya, V. \& Rhoades, B.E. 2014. Some
Convergence and Stability Results for the Kirk Multistep and Kirk-SP Fixed
Point Iterative Algorithms. Abstract and Applied Analysis
doi:10.1155/2014/806537.

\bibitem{DataF} G\"{u}rsoy, F., Karakaya, V. \& Rhoades, B.E. 2013. Data
dependence results of new multistep and S-iterative schemes for
contractive-like operators. Fixed Point Theory and Applications
doi:10.1186/1687-1812-2013-76.

\bibitem{Gursoy} G\"{u}rsoy, F., \& Karakaya, V. 2014. A Picard-S hybrid
type iteration method for solving a differential equation with retarded
argument. arXiv preprint arXiv:1403.2546.

\bibitem{Hammerlin} H\"{a}mmerlin, G. \& Hoffmann, K. H. 1991. Numerical
mathematics. Springer.

\bibitem{Hussain} Hussain, N., Rafiq, A., Damjanovi\'{c}, B. \& Lazovi\'{c},
R. 2011. On rate of convergence of various iterative schemes. Fixed Point
Theory and Applications 2011(1): 1-6.

\bibitem{Ishikawa} Ishikawa, S. 1974. Fixed points by a new iteration
method. Proceedings of the American Mathematical Society 44: 147-150.

\bibitem{Karahan} Karahan, I. \& Ozdemir, M. 2013. A general iterative
method for approximation of fixed points and their applications. Advances in
Fixed Point Theory 3(3).

\bibitem{K.kaya} Karakaya, V., G\"{u}rsoy, F. \& Ert\"{u}rk, M. 2014.
Comparison of The Speed of Convergence Among Various Iterative Schemes.
arXiv preprint arXiv:1402.6080.

\bibitem{Kkaya} Karakaya, V., G\"{u}rsoy, F., Do\u{g}an, K. \& Ert\"{u}rk,
M. 2103. Data Dependence Results for Multistep and CR Iterative Schemes in
the Class of Contractive-Like Operators. Abstract and Applied Analysis
doi:10.1155/2013/381980.

\bibitem{Karakaya} Karakaya, V., Do\u{g}an, K., G\"{u}rsoy, F. \& Ert\"{u}%
rk, M. 2013. Fixed Point of a New Three-Step Iteration Algorithm under
Contractive-Like Operators over Normed Spaces. Abstract and Applied Analysis
2013.

\bibitem{Khan} Khan, S. H. 2013. A Picard-Mann hybrid iterative process.
Fixed Point Theory and Applications 2013(1): 1-10.

\bibitem{Kirk} Kirk, W. 1971. On successive approximations for nonexpansive
mappings in Banach spaces. Glasgow Math. J. 12:6-9.

\bibitem{Krasnosel'skii} Krasnosel'skii, M.A. 1955. Two remarks on the
method of successive approximations. Uspekhi Matematicheskikh Nauk 10:
123-127.

\bibitem{Mann} Mann, W.R. 1953. Mean value methods in iteration. Proceedings
of the American Mathematical Society 4: 506-510.

\bibitem{Noor} Noor, M.A. 2000. New approximation schemes for general
variational inequalities. Journal of Mathematical Analysis and Applications
251: 217-229.

\bibitem{KM-I} Olatinwo, M. 2009. Some stability results for two hybrid
fixed point iterative algorithms in normed linear space. Matematique Vesnik
61: 247-256.

\bibitem{Ola} Olatinwo, M.O. 2008. Some stabilty results for two hybrid
fixed point iterative algorithms of Kirk-Ishikawa and Kirk-Mann type.
Journal of AdvancedMathematical Studies 1(1-2): 87-96.

\bibitem{Suantai} Phuengrattana, W. \& Suantai, S. 2012. Comparison of the
Rate of Convergence of Various Iterative Methods for the Class of Weak
Contractions in Banach Spaces. Thai Journal of Mathematics 11(1): 217-226.

\bibitem{SP} Phuengrattana, W. \& Suantai, S. 2011. On the rate of
convergence of Mann, Ishikawa, Noor and SP-iterations for continuous
functions on an arbitrary interval. Journal of Computational and Applied
Mathematics 235: 3006-3014.

\bibitem{Picard} Picard, E., (1890). "M\'{e}moire sur la th\'{e}orie des 
\'{e}quations aux d\'{e}riv\'{e}es partielles et la m\'{e}thode des
approximations successives", Journal de Math\'{e}matiques pures et appliqu%
\'{e}es, 6:145-210.

\bibitem{Popescu} Popescu, O. 2007. Picard iteration converges faster than
Mann iteration for a class of quasiontractive operators. Math. Commun. 12:
195-202.

\bibitem{Rhoades} Rhoades, B. E. \& \c{S}oltuz, S. M. 2004. The equivalence
between the convergences of Ishikawa and Mann iterations for an
asymptotically nonexpansive in the intermediate sense and strongly
successively pseudocontractive maps. Journal of mathematical analysis and
applications 289(1): 266-278.

\bibitem{Multistep} Rhoades, B. \& Soltuz, S.M. 2004. The equivalence
between Mann--Ishikawa iterations and multistep iteration. Nonlinear
Analysis: Theory, Methods \& Applications 58: 219-228.

\bibitem{Rhoades1} Rhoades, B. E. 1976. Comments on two fixed point
iteration methods. J. Math. Anal. Appl. 56(2): 741-750.

\bibitem{Sahu} Sahu, D. R. 2011. Applications of the S-iteration process to
constrained minimization problems and split feasibility problems. Fixed
Point Theory 12(1): 187-204.

\bibitem{SahuPet} Sahu, D. R. \& Petrusel, A. 2011. Strong convergence of
iterative methods by strictly pseudocontractive mappings in Banach spaces.
NonlinearAnalysis: Theory, Methods \&Applications 74(17): 6012-6023.

\bibitem{GenKrasnosel'skii} Schaefer, H. 1957. \"{U}ber die Methode
sukzessiver Approximationen. Jahresbericht der Deutschen
Mathematiker-Vereinigung 59: 131-140.

\bibitem{Data Is 2} Soltuz, S.M. \& Grosan, T. 2008. Data dependence for
Ishikawa iteration when dealing with contractive like operators. Fixed Point
Theory and Applications 2008: 242916(1-7).

\bibitem{SolOt1} Soltuz, S. M. \& Otrocol, D. 2007. Classical results via
Mann-Ishikawa iteration. Revue d'Analyse Numerique de l'Approximation 36(2).

\bibitem{Zxue} Sun, H. \& Xue, Z. 2013. Reconsiderations on the Equivalence
of Convergence between Mann and Ishikawa Iterations for Asymptotically
Pseudocontractive Mappings. Journal of Applied Mathematics 2013.

\bibitem{iam} Thianwan, S. 2009. Common fixed points of new iterations for
two asymptotically nonexpansive nonself-mappings in a Banach space. Journal
of Computational and Applied Mathematics 224: 688-695.

\bibitem{Weng} Weng, X. 1991. Fixed point iteration for local strictly
pseudocontractive mapping. Proc. Amer. Math. Soc. 113: 727-731.

\bibitem{ZXUE} Xue, Z. 2008. The comparison of the convergence speed between
Picard, Mann, Krasnoselskij and Ishikawa iterations in Banach spaces. Fixed
Point Theory Appl. 2008.

\bibitem{QR} Yuan, Q. \& Rhoades, B.E. 2008. Comments on the rate of
convergence between Mann and Ishikawa iterations applied to Zamfirescu
operators. Fixed Point Theory and Applications 2008.
\end{thebibliography}

\documentclass[global]{svjour}
\usepackage{amsfonts}
\usepackage{amsmath}

\input{tcilatex}
\begin{document}

\journalname{Name Your Journal}
\title{A Picard-S Iterative Scheme for Approximating Fixed Point of
Weak-Contraction Mappings}
\author{Faik G\"{U}RSOY}
\institute{Department of Mathematics, Yildiz Technical University, Davutpasa Campus,
Esenler, 34220 Istanbul, Turkey 
\email{faikgursoy02@hotmail.com}%
}
\offprints{Offprints Assistant}
\mail{Address for offprint requests}
\maketitle

\begin{abstract}
We study the convergence analysis of a Picard-S iteration method for a
particular class of weak-contraction mappings. Furthermore, we prove a data
dependence result for fixed point of the class of weak-contraction mappings
with the help of the Picard-S iteration methods.
\end{abstract}

\keywords{Picard-S iterative scheme -- Weak-Contraction mappings --
Convergence -- Rate of Convergence- Data Dependence }

\section{Introduction}

Most of the problems that arise in different disciplines of science can be
formulated by the equations in the form%
\begin{equation}
Fx=0\text{,}  \label{eqn1}
\end{equation}%
where $F$ is some function. The equations given by (1) can easily be
reformulated as the fixed point equations of type 
\begin{equation}
Tx=x\text{.}  \label{eqn2}
\end{equation}%
where $T$ is a self-map of an ambient space $X$ and $x\in X$. These
equations are often classified as linear or nonlinear, depending on whether
the mappings used in the equation is linear with respect to the variables.
Over the years, a considerable attention has been paid to solving such
equations by using different techniques such as direct and iterative
methods. In case of linear equations, both direct and iterative methods are
used to obtain solutions of the equations. But in case of nonlinear
equations, due to various reasons, direct methods can be impractical or fail
in solving equations, and thus iterative methods become a viable
alternative. Nonlinear problems are of importance interest to
mathematicians, physicists and engineers and many other scientists, simply
because most systems are intrinsically nonlinear in nature. That is why,
researchers in various disciplines of sciences are often faced with the
solving such problems. It would be hard to fudge that the role of iterative
approximation of fixed points have played in the recent progress of
nonlinear science. Indeed, the phrase iterative approximation has been
introduced to describe repetition-based researches into nonlinear problems
that are inaccessible to analytic methods. For this reason, the iterative
approximation of fixed points has become one of the major and basic tools in
the theory of equations, and as a result, numerous iterative methods have
been introduced or improved and studied for many years in detail from
various points of aspects by a wide audience of researchers, see, [1-14,
16-42, 44, 45].

In this paper, we show that a Picard-S iteration method \cite{Gursoy} can be
used to approximate fixed point of weak-contraction mappings. Also, we show
that this iteration method is equivalent and converges faster than CR
iteration method \cite{CR} for the aforementioned class of mappings.
Furthermore, by providing an example, it is shown that the Picard-S
iteration method converges faster than CR iteration method and hence also
faster than all Picard \cite{Picard}, Mann \cite{Mann}, Ishikawa \cite%
{Ishikawa}, Noor \cite{Noor}, SP \cite{SP}, S \cite{S} and some other
iteration methods in the existing literature when applied to
weak-contraction mappings. Finally, a data dependence result is proven for
fixed point of weak-contraction mappings with the help of the Picard-S
iteration method.

Throughout this paper the set of all positive integers and zero is shown by $%
\mathbb{N}
$. Let $B$ be a Banach space, $D$ be a nonempty closed convex subset of $B$
and $T$ a self-map of $D$. An element $x_{\ast }$ of $D$ is called a fixed
point of $T$ if and only if $Tx_{\ast }=x_{\ast }$. The set of all fixed
point of $T$ denoted by $F_{T}$. Let $\left\{ a_{n}^{i}\right\}
_{n=0}^{\infty }$, $i\in \left\{ 0,1,2\right\} $ be real sequences in $\left[
0,1\right] $ satisfying certain control condition(s).

Renowned Picard iteration method \cite{Picard} is formulated as follow%
\begin{equation}
\left\{ 
\begin{array}{c}
p_{0}\in D\text{, \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ } \\ 
p_{n+1}=Tp_{n}\text{, }n\in 
\mathbb{N}
\text{,}%
\end{array}%
\right.   \label{eqn3}
\end{equation}%
and generally used to approximate fixed points of contraction mappings
satisfying: for all $x$, $y\in B$ there exists a $\delta \in \left(
0,1\right) $ such that%
\begin{equation}
\left\Vert Tx-Ty\right\Vert \leq \delta \left\Vert x-y\right\Vert \text{.}
\label{eqn4}
\end{equation}%
The following iteration methods are known as Noor \cite{Noor} and SP \cite%
{SP} iterations, respectively:%
\begin{equation}
\left\{ 
\begin{array}{c}
\omega _{0}\in D\text{, \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \
\ \ \ \ \ } \\ 
\omega _{n+1}=\left( 1-a_{n}^{0}\right) \omega _{n}+a_{n}^{0}T\varpi _{n}%
\text{, \ \ } \\ 
\varpi _{n}=\left( 1-a_{n}^{1}\right) \omega _{n}+a_{n}^{1}T\rho _{n}\text{, 
} \\ 
\rho _{n}=\left( 1-a_{n}^{2}\right) \omega _{n}+a_{n}^{2}T\omega _{n}\text{, 
}n\in 
\mathbb{N}
\text{,}%
\end{array}%
\right.   \label{eqn5}
\end{equation}%
\begin{equation}
\left\{ 
\begin{array}{c}
q_{0}\in D\text{,\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \
\ } \\ 
q_{n+1}=\left( 1-a_{n}^{0}\right) r_{n}+a_{n}^{0}Tr_{n}\text{, \ \ } \\ 
r_{n}=\left( 1-a_{n}^{1}\right) s_{n}+a_{n}^{1}Ts_{n}\text{,} \\ 
s_{n}=\left( 1-a_{n}^{2}\right) q_{n}+a_{n}^{2}Tq_{n}\text{, }n\in 
\mathbb{N}
\text{,}%
\end{array}%
\right.   \label{eqn6}
\end{equation}

\begin{remark}
(i) If $a_{n}^{2}=0$ for each $n\in 
\mathbb{N}
$ , then the Noor iteration method reduces to iterative method of Ishikawa 
\cite{Ishikawa}.

(ii) If $a_{n}^{2}=0$ for each $n\in 
\mathbb{N}
$ , then the SP iteration method reduces to iterative method of Thianwan 
\cite{iam}. 

(iii) When $a_{n}^{1}=$ $a_{n}^{2}=0$ for each $n\in 
\mathbb{N}
$, then both Noor and SP iteration methods reduce to an iteration method due
to Mann \cite{Mann}.  
\end{remark}

Recently, G\"{u}rsoy and Karakaya \cite{Gursoy} introduced a Picard-S
iterative scheme as follows: 
\begin{equation}
\left\{ 
\begin{array}{c}
x_{0}\in D\text{, \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \
\ \ \ \ \ \ \ \ \ \ \ } \\ 
x_{n+1}=Ty_{n}\text{, \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \
\ \ \ \ \ \ \ } \\ 
y_{n}=\left( 1-a_{n}^{1}\right) Tx_{n}+a_{n}^{1}Tz_{n}\text{, \ \ \ \ \ \ }
\\ 
z_{n}=\left( 1-a_{n}^{2}\right) x_{n}+a_{n}^{2}Tx_{n}\text{, }n\in 
\mathbb{N}
\text{,}%
\end{array}%
\right.   \label{eqn7}
\end{equation}%
The following definitions and lemmas will be needed in obtaining the main
results of this article.

\begin{definition}
\cite{Berinde} Let $\left\{ a_{n}\right\} _{n=0}^{\infty }$ and $\left\{
b_{n}\right\} _{n=0}^{\infty }$ be two sequences of real numbers with limits 
$a$ and $b$, respectively. Assume that there exists%
\begin{equation}
\underset{n\rightarrow \infty }{\lim }\frac{\left\vert a_{n}-a\right\vert }{%
\left\vert b_{n}-b\right\vert }=l\text{.}  \label{eqn8}
\end{equation}

(i) If $l=0$, the we say that $\left\{ a_{n}\right\} _{n=0}^{\infty }$
converges faster to $a$ than $\left\{ b_{n}\right\} _{n=0}^{\infty }$ to $b$.

(ii) If $0<l<\infty $, then we say that $\left\{ a_{n}\right\}
_{n=0}^{\infty }$ and $\left\{ b_{n}\right\} _{n=0}^{\infty }$ have the same
rate of convergence.
\end{definition}

\begin{definition}
\cite{Berinde} Assume that for two fixed point iteration processes $\left\{
u_{n}\right\} _{n=0}^{\infty }$ and $\left\{ v_{n}\right\} _{n=0}^{\infty }$
both converging to the same fixed point $p$, the following error predictions%
\begin{equation}
\left\Vert u_{n}-p\right\Vert \leq a_{n}\text{ for all }n\in 
\mathbb{N}
\text{,}  \label{eqn9}
\end{equation}%
\begin{equation}
\left\Vert v_{n}-p\right\Vert \leq b_{n}\text{ for all }n\in 
\mathbb{N}
\text{,}  \label{eqn10}
\end{equation}%
are available where $\left\{ a_{n}\right\} _{n=0}^{\infty }$ and $\left\{
b_{n}\right\} _{n=0}^{\infty }$ are two sequences of positive numbers
(converging to zero). If $\left\{ a_{n}\right\} _{n=0}^{\infty }$ converges
faster than $\left\{ b_{n}\right\} _{n=0}^{\infty }$, then $\left\{
u_{n}\right\} _{n=0}^{\infty }$ converges faster than $\left\{ v_{n}\right\}
_{n=0}^{\infty }$ to $p$.
\end{definition}

\begin{definition}
\cite{Vasile} Let $\left( B,\left\Vert \cdot \right\Vert \right) $ be a
Banach space. A map $T:B\rightarrow B$ is called weak-contraction if there
exist a constant $\delta \in \left( 0,1\right) $ and some $L\geq 0$ such that%
\begin{equation}
\left\Vert Tx-Ty\right\Vert \leq \delta \left\Vert x-y\right\Vert
+L\left\Vert y-Ty\right\Vert \text{, for all }x\text{, }y\in B\text{.}
\label{eqn11}
\end{equation}
\end{definition}

\begin{definition}
\cite{Vasile} Let $T$,$\widetilde{T}:B\rightarrow B$ be two operators. We
say that $\widetilde{T}$ is an approximate operator of $T$ if for all $x\in B
$ and for a fixed $\varepsilon >0$ we have 
\begin{equation}
\left\Vert Tx-\widetilde{T}x\right\Vert \leq \varepsilon .  \label{eqn12}
\end{equation}
\end{definition}

\begin{lemma}
\cite{Weng}Let $\left\{ \beta _{n}\right\} _{n=0}^{\infty }$ and $\left\{
\rho _{n}\right\} _{n=0}^{\infty }$ be nonnegative real sequences satisfying
the following inequality:%
\begin{equation}
\beta _{n+1}\leq \left( 1-\lambda _{n}\right) \beta _{n}+\rho _{n}\text{,}
\label{eqn13}
\end{equation}%
where $\lambda _{n}\in \left( 0,1\right) $, for all $n\geq n_{0}$, $%
\dsum\nolimits_{n=1}^{\infty }\lambda _{n}=\infty $, and $\frac{\rho _{n}}{%
\lambda _{n}}\rightarrow 0$ as $n\rightarrow \infty $. Then $%
\lim_{n\rightarrow \infty }\beta _{n}=0$.
\end{lemma}

\begin{lemma}
\cite{Data Is 2} Let $\left\{ \beta _{n}\right\} _{n=0}^{\infty }$ be a
nonnegative sequence for which one assumes there exists $n_{0}\in 
\mathbb{N}
$, such that for all $n\geq n_{0}$ one has satisfied the inequality 
\begin{equation}
\beta _{n+1}\leq \left( 1-\mu _{n}\right) \beta _{n}+\mu _{n}\gamma _{n}%
\text{,}  \label{eqn15}
\end{equation}%
where $\mu _{n}\in \left( 0,1\right) ,$ for all $n\in 
\mathbb{N}
$, $\sum\limits_{n=0}^{\infty }\mu _{n}=\infty $ and $\gamma _{n}\geq 0$, $%
\forall n\in 
\mathbb{N}
$. Then the following inequality holds 
\begin{equation}
0\leq \lim \sup_{n\rightarrow \infty }\beta _{n}\leq \lim \sup_{n\rightarrow
\infty }\gamma _{n}.  \label{eqn16}
\end{equation}
\end{lemma}

\section{Main Results}

\begin{theorem}
Let $T:D\rightarrow D$ be a weak-contraction map satisfying condition (11)
with $F_{T}\neq \emptyset $ and $\left\{ x_{n}\right\} _{n=0}^{\infty }$ an
iterative sequence defined by (7) with real sequences $\left\{
a_{n}^{i}\right\} _{n=0}^{\infty }$, $i\in \left\{ 1,2\right\} $ in $\left[
0,1\right] $ satisfying $\sum_{k=0}^{\infty }a_{k}^{1}a_{k}^{2}=\infty $.
Then $\ \left\{ x_{n}\right\} _{n=0}^{\infty }$ converges to a unique fixed
point $u^{\ast }$of $T$.
\end{theorem}

\begin{proof}
Uniqueness of $u^{\ast }$ comes from condition (11). Using Picard-S
iterative scheme (7) and condition (11), we obtain%
\begin{eqnarray}
\left\Vert z_{n}-u^{\ast }\right\Vert  &\leq &\left( 1-a_{n}^{2}\right)
\left\Vert x_{n}-u^{\ast }\right\Vert +a_{n}^{2}\left\Vert Tx_{n}-Tu^{\ast
}\right\Vert   \notag \\
&\leq &\left( 1-a_{n}^{2}\right) \left\Vert x_{n}-u^{\ast }\right\Vert
+a_{n}^{2}\delta \left\Vert x_{n}-u^{\ast }\right\Vert +a_{n}^{2}L\left\Vert
u^{\ast }-Tu^{\ast }\right\Vert   \notag \\
&=&\left[ 1-a_{n}^{2}\left( 1-\delta \right) \right] \left\Vert
x_{n}-u^{\ast }\right\Vert \text{,}  \label{eqn17}
\end{eqnarray}%
\begin{eqnarray}
\left\Vert y_{n}-u^{\ast }\right\Vert  &\leq &\left( 1-a_{n}^{1}\right)
\left\Vert Tx_{n}-Tu^{\ast }\right\Vert +a_{n}^{1}\left\Vert Tz_{n}-Tu^{\ast
}\right\Vert   \notag \\
&\leq &\left( 1-a_{n}^{1}\right) \delta \left\Vert x_{n}-u^{\ast
}\right\Vert +a_{n}^{1}\delta \left\Vert z_{n}-u^{\ast }\right\Vert \text{,}
\label{eqn18}
\end{eqnarray}%
\begin{equation}
\left\Vert x_{n+1}-u^{\ast }\right\Vert \leq \delta \left\Vert y_{n}-u^{\ast
}\right\Vert \text{.}  \label{eqn19}
\end{equation}%
Combining (16), (17) and (18)%
\begin{equation}
\left\Vert x_{n+1}-u^{\ast }\right\Vert \leq \delta ^{2}\left[
1-a_{n}^{1}a_{n}^{2}\left( 1-\delta \right) \right] \left\Vert x_{n}-u^{\ast
}\right\Vert \text{.}  \label{eqn20}
\end{equation}%
By induction%
\begin{eqnarray}
\left\Vert x_{n+1}-u^{\ast }\right\Vert  &\leq &\delta ^{2\left( n+1\right)
}\dprod\nolimits_{k=0}^{n}\left[ 1-a_{k}^{1}a_{k}^{2}\left( 1-\delta \right) %
\right] \left\Vert x_{0}-u^{\ast }\right\Vert   \notag \\
&\leq &\delta ^{2\left( n+1\right) }\left\Vert x_{0}-u^{\ast }\right\Vert
^{n+1}e^{-\left( 1-\delta \right) \dsum\nolimits_{k=0}^{n}a_{k}^{1}a_{k}^{2}}%
\text{.}  \label{eqn21}
\end{eqnarray}%
Since $\sum_{k=0}^{\infty }a_{k}^{1}a_{k}^{2}=\infty $,%
\begin{equation}
e^{-\left( 1-\delta \right)
\dsum\nolimits_{k=0}^{n}a_{k}^{1}a_{k}^{2}}\rightarrow 0\text{ as }%
n\rightarrow \infty \text{,}  \label{eqn22}
\end{equation}%
which implies $\lim_{n\rightarrow \infty }\left\Vert x_{n}-u^{\ast
}\right\Vert $.
\end{proof}

\begin{theorem}
Let $T:D\rightarrow D$ with fixed point $u^{\ast }\in F_{T}\neq \emptyset $
be as in Theorem 1 and $\{q_{n}\}_{n=0}^{\infty }$, $\{x_{n}\}_{n=0}^{\infty
}$ two iterative sequences defined by SP (6) and Picard-S (7) iteration
methods with real sequences $\left\{ a_{n}^{i}\right\} _{n=0}^{\infty }$, $%
i\in \left\{ 0,1,2\right\} $ in $\left[ 0,1\right] $ satisfying $%
\sum_{k=0}^{n}a_{k}^{1}a_{k}^{2}=\infty $. Then the following are equivalent:

(i) $\lim_{n\rightarrow \infty }\left\Vert x_{n}-u^{\ast }\right\Vert =0$;

(ii) $\lim_{n\rightarrow \infty }\left\Vert q_{n}-u^{\ast }\right\Vert =0$.

\begin{proof}
(i)$\Rightarrow $(ii): It follows from (6), (7), and condition (11) that%
\begin{eqnarray}
\left\Vert x_{n+1}-q_{n+1}\right\Vert  &=&\left\Vert \left(
1-a_{n}^{0}\right) \left( Ty_{n}-r_{n}\right) +a_{n}^{0}\left(
Ty_{n}-Tr_{n}\right) \right\Vert   \label{eqn23} \\
&\leq &\left( 1-a_{n}^{0}\right) \left\Vert Ty_{n}-r_{n}\right\Vert
+a_{n}^{0}\left\Vert Ty_{n}-Tr_{n}\right\Vert   \notag \\
&\leq &\left[ 1-a_{n}^{0}\left( 1-\delta \right) \right] \left\Vert
y_{n}-r_{n}\right\Vert +\left[ 1-a_{n}^{0}\left( 1-L\right) \right]
\left\Vert y_{n}-Ty_{n}\right\Vert \text{,}  \notag
\end{eqnarray}%
\begin{eqnarray}
\left\Vert y_{n}-r_{n}\right\Vert  &=&\left\Vert \left( 1-a_{n}^{1}\right)
\left( Tx_{n}-s_{n}\right) +a_{n}^{1}\left( Tz_{n}-Ts_{n}\right) \right\Vert 
\label{eqn24} \\
&\leq &\left( 1-a_{n}^{1}\right) \left\Vert Tx_{n}-s_{n}\right\Vert
+a_{n}^{1}\left\Vert Tz_{n}-Ts_{n}\right\Vert   \notag \\
&\leq &\left( 1-a_{n}^{1}\right) \left\Vert Tx_{n}-s_{n}\right\Vert
+a_{n}^{1}\delta \left\Vert z_{n}-s_{n}\right\Vert +a_{n}^{1}L\left\Vert
z_{n}-Tz_{n}\right\Vert \text{,}  \notag
\end{eqnarray}%
\begin{eqnarray}
\left\Vert Tx_{n}-s_{n}\right\Vert  &=&\left\Vert \left( 1-a_{n}^{2}\right)
\left( Tx_{n}-q_{n}\right) +a_{n}^{2}\left( Tx_{n}-Tq_{n}\right) \right\Vert 
\label{eqn25} \\
&\leq &\left[ 1-a_{n}^{2}\left( 1-\delta \right) \right] \left\Vert
x_{n}-q_{n}\right\Vert +\left[ 1-a_{n}^{2}\left( 1-L\right) \right]
\left\Vert x_{n}-Tx_{n}\right\Vert \text{,}  \notag
\end{eqnarray}%
\begin{eqnarray}
\left\Vert z_{n}-s_{n}\right\Vert  &\leq &\left( 1-a_{n}^{2}\right)
\left\Vert x_{n}-q_{n}\right\Vert +a_{n}^{2}\left\Vert
Tx_{n}-Tq_{n}\right\Vert   \label{eqn26} \\
&\leq &\left[ 1-a_{n}^{2}\left( 1-\delta \right) \right] \left\Vert
x_{n}-q_{n}\right\Vert +a_{n}^{2}L\left\Vert x_{n}-Tx_{n}\right\Vert \text{.}
\notag
\end{eqnarray}%
Combining (22), (23), (24), and (25)%
\begin{eqnarray}
\left\Vert x_{n+1}-q_{n+1}\right\Vert  &\leq &\left[ 1-a_{n}^{0}\left(
1-\delta \right) \right] \left[ 1-a_{n}^{1}\left( 1-\delta \right) \right] %
\left[ 1-a_{n}^{2}\left( 1-\delta \right) \right] \left\Vert
x_{n}-q_{n}\right\Vert   \label{eqn27} \\
&&+\left[ 1-a_{n}^{0}\left( 1-\delta \right) \right] \left\{ \left(
1-a_{n}^{1}\right) \left[ 1-a_{n}^{2}\left( 1-L\right) \right]
+a_{n}^{1}a_{n}^{2}\delta L\right\} \left\Vert x_{n}-Tx_{n}\right\Vert  
\notag \\
&&+\left[ 1-a_{n}^{0}\left( 1-\delta \right) \right] a_{n}^{1}L\left\Vert
z_{n}-Tz_{n}\right\Vert +\left[ 1-a_{n}^{0}\left( 1-L\right) \right]
\left\Vert y_{n}-Ty_{n}\right\Vert \text{.}  \notag
\end{eqnarray}%
It follows from the facts $\delta \in \left( 0,1\right) $ and $a_{n}^{i}\in %
\left[ 0,1\right] $, $\forall n\in 
\mathbb{N}
$, $i\in \left\{ 0,1,2\right\} $ that%
\begin{equation}
\left[ 1-a_{n}^{0}\left( 1-\delta \right) \right] \left[ 1-a_{n}^{1}\left(
1-\delta \right) \right] \left[ 1-a_{n}^{2}\left( 1-\delta \right) \right]
<1-a_{n}^{1}a_{n}^{2}\left( 1-\delta \right) \text{.}  \label{eqn28}
\end{equation}%
Hence, inequality (26) becomes%
\begin{eqnarray}
\left\Vert x_{n+1}-q_{n+1}\right\Vert  &\leq &\left[ 1-a_{n}^{1}a_{n}^{2}%
\left( 1-\delta \right) \right] \left\Vert x_{n}-q_{n}\right\Vert 
\label{eqn29} \\
&&+\left[ 1-a_{n}^{0}\left( 1-\delta \right) \right] \left\{ \left(
1-a_{n}^{1}\right) \left[ 1-a_{n}^{2}\left( 1-L\right) \right]
+a_{n}^{1}a_{n}^{2}\delta L\right\} \left\Vert x_{n}-Tx_{n}\right\Vert  
\notag \\
&&+\left[ 1-a_{n}^{0}\left( 1-\delta \right) \right] a_{n}^{1}L\left\Vert
z_{n}-Tz_{n}\right\Vert +\left[ 1-a_{n}^{0}\left( 1-L\right) \right]
\left\Vert y_{n}-Ty_{n}\right\Vert \text{.}  \notag
\end{eqnarray}%
Denote that%
\begin{eqnarray}
\beta _{n} &:&=\left\Vert x_{n}-q_{n}\right\Vert \text{, \ \ }  \label{eqn30}
\\
\lambda _{n} &:&=a_{n}^{1}a_{n}^{2}\left( 1-\delta \right) \in \left(
0,1\right) \text{, \ }  \notag \\
\rho _{n} &:&=\left[ 1-a_{n}^{0}\left( 1-\delta \right) \right] \left\{
\left( 1-a_{n}^{1}\right) \left[ 1-a_{n}^{2}\left( 1-L\right) \right]
+a_{n}^{1}a_{n}^{2}\delta L\right\} \left\Vert x_{n}-Tx_{n}\right\Vert  
\notag \\
&&+\left[ 1-a_{n}^{0}\left( 1-\delta \right) \right] a_{n}^{1}L\left\Vert
z_{n}-Tz_{n}\right\Vert +\left[ 1-a_{n}^{0}\left( 1-L\right) \right]
\left\Vert y_{n}-Ty_{n}\right\Vert \text{.}  \notag
\end{eqnarray}%
Since $\lim_{n\rightarrow \infty }\left\Vert x_{n}-u^{\ast }\right\Vert =0$
and $Tu^{\ast }=u^{\ast }$%
\begin{equation}
\lim_{n\rightarrow \infty }\left\Vert x_{n}-Tx_{n}\right\Vert
=\lim_{n\rightarrow \infty }\left\Vert y_{n}-Ty_{n}\right\Vert
=\lim_{n\rightarrow \infty }\left\Vert z_{n}-Tz_{n}\right\Vert =0\text{,}
\label{eqn31}
\end{equation}%
which implies $\frac{\rho _{n}}{\lambda _{n}}\rightarrow 0$ as $n\rightarrow
\infty $. Therefore, inequality (28) perform all assumptions in Lemma 1 and
thus we obtain $\lim_{n\rightarrow \infty }\left\Vert x_{n}-q_{n}\right\Vert
=0$. Since%
\begin{equation}
\left\Vert q_{n}-u^{\ast }\right\Vert \leq \left\Vert x_{n}-q_{n}\right\Vert
+\left\Vert x_{n}-u^{\ast }\right\Vert \rightarrow 0\text{ as }n\rightarrow
\infty \text{,}  \label{eqn32}
\end{equation}%
$\lim_{n\rightarrow \infty }\left\Vert q_{n}-u^{\ast }\right\Vert =0$.

(ii)$\Rightarrow $(i): It follows from (6), (7), and condition (11) that%
\begin{eqnarray}
\left\Vert q_{n+1}-x_{n+1}\right\Vert  &=&\left\Vert
r_{n}-Ty_{n}+a_{n}^{0}\left( Tr_{n}-r_{n}\right) \right\Vert   \label{eqn33}
\\
&\leq &\delta \left\Vert r_{n}-y_{n}\right\Vert +\left( 1+a_{n}^{0}+L\right)
\left\Vert r_{n}-Tr_{n}\right\Vert \text{,}  \notag
\end{eqnarray}%
\begin{eqnarray}
\left\Vert r_{n}-y_{n}\right\Vert  &\leq &\left( 1-a_{n}^{1}\right)
\left\Vert s_{n}-Tx_{n}\right\Vert +a_{n}^{1}\left\Vert
Ts_{n}-Tz_{n}\right\Vert   \label{eqn34} \\
&\leq &\left( 1-a_{n}^{1}\right) \left\Vert s_{n}-Tx_{n}\right\Vert
+a_{n}^{1}\delta \left\Vert s_{n}-z_{n}\right\Vert +a_{n}^{1}L\left\Vert
s_{n}-Ts_{n}\right\Vert \text{,}  \notag
\end{eqnarray}%
\begin{eqnarray}
\left\Vert s_{n}-Tx_{n}\right\Vert  &\leq &\left\Vert
Ts_{n}-Tx_{n}\right\Vert +\left\Vert s_{n}-Ts_{n}\right\Vert   \label{eqn35}
\\
&\leq &\delta \left\Vert s_{n}-x_{n}\right\Vert +\left( 1+L\right)
\left\Vert s_{n}-Ts_{n}\right\Vert   \notag \\
&\leq &\delta \left\Vert q_{n}-x_{n}\right\Vert +\delta a_{n}^{2}\left\Vert
Tq_{n}-q_{n}\right\Vert +\left( 1+L\right) \left\Vert
s_{n}-Ts_{n}\right\Vert \text{,}  \notag
\end{eqnarray}%
\begin{eqnarray}
\left\Vert s_{n}-z_{n}\right\Vert  &\leq &\left( 1-a_{n}^{2}\right)
\left\Vert q_{n}-x_{n}\right\Vert +a_{n}^{2}\left\Vert
Tq_{n}-Tx_{n}\right\Vert   \label{eqn36} \\
&\leq &\left[ 1-a_{n}^{2}\left( 1-\delta \right) \right] \left\Vert
q_{n}-x_{n}\right\Vert +a_{n}^{2}L\left\Vert q_{n}-Tq_{n}\right\Vert \text{.}
\notag
\end{eqnarray}%
Combining (32), (33), (34), and (35)%
\begin{eqnarray}
\left\Vert q_{n+1}-x_{n+1}\right\Vert  &\leq &\delta ^{2}\left[
1-a_{n}^{1}a_{n}^{2}\left( 1-\delta \right) \right] \left\Vert
q_{n}-x_{n}\right\Vert   \label{eqn37} \\
&&+\delta ^{2}a_{n}^{2}\left[ 1-a_{n}^{1}\left( 1-L\right) \right]
\left\Vert q_{n}-Tq_{n}\right\Vert   \notag \\
&&+\left( 1+a_{n}^{0}+L\right) \left\Vert r_{n}-Tr_{n}\right\Vert +\delta
\left( 1-a_{n}^{1}+L\right) \left\Vert s_{n}-Ts_{n}\right\Vert \text{.} 
\notag
\end{eqnarray}%
Since $\delta \in \left( 0,1\right) $%
\begin{equation}
\delta ^{2}\left[ 1-a_{n}^{1}a_{n}^{2}\left( 1-\delta \right) \right]
<1-a_{n}^{1}a_{n}^{2}\left( 1-\delta \right) \text{.}  \label{eqn38}
\end{equation}%
Hence, inequality (36) becomes%
\begin{eqnarray}
\left\Vert q_{n+1}-x_{n+1}\right\Vert  &\leq &\left[ 1-a_{n}^{1}a_{n}^{2}%
\left( 1-\delta \right) \right] \left\Vert q_{n}-x_{n}\right\Vert 
\label{eqn39} \\
&&+\delta ^{2}a_{n}^{2}\left[ 1-a_{n}^{1}\left( 1-L\right) \right]
\left\Vert q_{n}-Tq_{n}\right\Vert   \notag \\
&&+\left( 1+a_{n}^{0}+L\right) \left\Vert r_{n}-Tr_{n}\right\Vert +\delta
\left( 1-a_{n}^{1}+L\right) \left\Vert s_{n}-Ts_{n}\right\Vert \text{.} 
\notag
\end{eqnarray}%
Denote that%
\begin{eqnarray}
\beta _{n} &:&=\left\Vert q_{n}-x_{n}\right\Vert \text{, \ \ }  \label{eqn40}
\\
\lambda _{n} &:&=a_{n}^{1}a_{n}^{2}\left( 1-\delta \right) \in \left(
0,1\right) \text{, \ }  \notag \\
\rho _{n} &:&=\delta ^{2}a_{n}^{2}\left[ 1-a_{n}^{1}\left( 1-L\right) \right]
\left\Vert q_{n}-Tq_{n}\right\Vert   \notag \\
&&+\left( 1+a_{n}^{0}+L\right) \left\Vert r_{n}-Tr_{n}\right\Vert +\delta
\left( 1-a_{n}^{1}+L\right) \left\Vert s_{n}-Ts_{n}\right\Vert \text{.} 
\notag
\end{eqnarray}%
Since $\lim_{n\rightarrow \infty }\left\Vert q_{n}-u^{\ast }\right\Vert =0$
and $Tu^{\ast }=u^{\ast }$%
\begin{equation}
\lim_{n\rightarrow \infty }\left\Vert q_{n}-Tq_{n}\right\Vert
=\lim_{n\rightarrow \infty }\left\Vert r_{n}-Tr_{n}\right\Vert
=\lim_{n\rightarrow \infty }\left\Vert s_{n}-Ts_{n}\right\Vert =0\text{,}
\label{eqn41}
\end{equation}%
which implies $\frac{\rho _{n}}{\lambda _{n}}\rightarrow 0$ as $n\rightarrow
\infty $. Therefore, inequality (38) perform all assumptions in Lemma 1 and
thus we obtain $\lim_{n\rightarrow \infty }\left\Vert q_{n}-x_{n}\right\Vert
=0$. Since%
\begin{equation}
\left\Vert x_{n}-u^{\ast }\right\Vert \leq \left\Vert q_{n}-x_{n}\right\Vert
+\left\Vert q_{n}-u^{\ast }\right\Vert \rightarrow 0\text{ as }n\rightarrow
\infty \text{,}  \label{eqn42}
\end{equation}%
$\lim_{n\rightarrow \infty }\left\Vert x_{n}-u^{\ast }\right\Vert =0$.
\end{proof}
\end{theorem}

Taking R. Chugh et al.'s result (\cite{CR}, Corollary 3.2) into account,
Theorem 2 leads to the following corollary under weaker assumption:

\begin{corollary}
Let $T:D\rightarrow D$ with fixed point $u^{\ast }\in F_{T}\neq \emptyset $
be as in Theorem 1. Then the followings are equivalent:

1)The Picard iteration method (3) converges to $u^{\ast }$,

2) The Mann iteration method \cite{Mann} converges to $u^{\ast }$,

3) The Ishikawa iteration method \cite{Ishikawa} converges to $u^{\ast }$,

4) The Noor iteration method (5) converges to $u^{\ast }$,

5) S-iteration method \cite{S} converges to $u^{\ast }$,

6) The SP-iteration method (6) converges to $u^{\ast }$,

7) CR-iteration method \cite{CR} converges to $u^{\ast }$,

8) The Picard-S iteration method (7) converges to $u^{\ast }$.
\end{corollary}

\begin{theorem}
Let $T:D\rightarrow D$ with fixed point $u^{\ast }\in F_{T}\neq \emptyset $
be as in Theorem 1. Suppose that $\left\{ \omega _{n}\right\} _{n=0}^{\infty
}$, $\left\{ q_{n}\right\} _{n=0}^{\infty }$ and $\left\{ x_{n}\right\}
_{n=0}^{\infty }$ are iterative sequences, respectively, defined by Noor
(5), SP (6) and Picard-S (7) iterative schemes with real sequences $\left\{
a_{n}^{i}\right\} _{n=0}^{\infty }\subset \left[ 0,1\right] $, $i\in \left\{
0\text{,}1\text{,}2\right\} $ satisfying

(i) $0\leq a_{n}^{i}<\frac{1}{1+\delta }$,

(ii) $\lim_{n\rightarrow \infty }a_{n}^{i}=0$.

Then the iterative sequence defined by (7) converges faster than the
iterative sequences defined by (5) and (6) to a unique fixed point of $T$,
provided that the initial point is the same for all iterations.
\end{theorem}

\begin{proof}
From inequality (20), we have%
\begin{equation}
\left\Vert x_{n+1}-u^{\ast }\right\Vert \leq \delta ^{2\left( n+1\right)
}\left\Vert x_{0}-u^{\ast }\right\Vert ^{n+1}\dprod\nolimits_{k=0}^{n}\left[
1-a_{k}^{1}a_{k}^{2}\left( 1-\delta \right) \right]   \label{eqn43}
\end{equation}%
Using (6) we obtain%
\begin{eqnarray}
\left\Vert q_{n+1}-u^{\ast }\right\Vert  &=&\left\Vert \left(
1-a_{n}^{0}\right) r_{n}+a_{n}^{0}Tr_{n}-u^{\ast }\right\Vert   \label{eqn44}
\\
&\geq &\left( 1-a_{n}^{0}\right) \left\Vert r_{n}-u^{\ast }\right\Vert
-a_{n}^{0}\left\Vert Tr_{n}-Tu^{\ast }\right\Vert   \notag \\
&\geq &\left[ 1-a_{n}^{0}\left( 1+\delta \right) \right] \left\Vert
r_{n}-u^{\ast }\right\Vert   \notag \\
&\geq &\left[ 1-a_{n}^{0}\left( 1+\delta \right) \right] \left\{ \left(
1-a_{n}^{1}\right) \left\Vert s_{n}-u^{\ast }\right\Vert -a_{n}^{1}\delta
\left\Vert s_{n}-u^{\ast }\right\Vert \right\}   \notag \\
&=&\left[ 1-a_{n}^{0}\left( 1+\delta \right) \right] \left[
1-a_{n}^{1}\left( 1+\delta \right) \right] \left\Vert s_{n}-u^{\ast
}\right\Vert   \notag \\
&\geq &\left[ 1-a_{n}^{0}\left( 1+\delta \right) \right] \left[
1-a_{n}^{1}\left( 1+\delta \right) \right] \left\{ \left( 1-a_{n}^{2}\right)
\left\Vert q_{n}-u^{\ast }\right\Vert -a_{n}^{2}\delta \left\Vert
q_{n}-u^{\ast }\right\Vert \right\}   \notag \\
&=&\left[ 1-a_{n}^{0}\left( 1+\delta \right) \right] \left[
1-a_{n}^{1}\left( 1+\delta \right) \right] \left[ 1-a_{n}^{2}\left( 1+\delta
\right) \right] \left\Vert q_{n}-u^{\ast }\right\Vert   \notag \\
&\geq &\cdots   \notag \\
&\geq &\left\Vert q_{0}-u^{\ast }\right\Vert ^{n+1}\prod\limits_{k=0}^{n}
\left[ 1-a_{k}^{0}\left( 1+\delta \right) \right] \left[ 1-a_{k}^{1}\left(
1+\delta \right) \right] \left[ 1-a_{k}^{2}\left( 1+\delta \right) \right] 
\text{.}  \notag
\end{eqnarray}%
Using now (42) and (43)%
\begin{equation}
\frac{\left\Vert x_{n+1}-u^{\ast }\right\Vert }{\left\Vert q_{n+1}-u^{\ast
}\right\Vert }\leq \frac{\delta ^{2\left( n+1\right) }\left\Vert
x_{0}-u^{\ast }\right\Vert ^{n+1}\dprod\nolimits_{k=0}^{n}\left[
1-a_{k}^{1}a_{k}^{2}\left( 1-\delta \right) \right] }{\left\Vert
q_{0}-u^{\ast }\right\Vert ^{n+1}\prod\limits_{k=0}^{n}\left[
1-a_{k}^{0}\left( 1+\delta \right) \right] \left[ 1-a_{k}^{1}\left( 1+\delta
\right) \right] \left[ 1-a_{k}^{2}\left( 1+\delta \right) \right] }\text{.}
\label{eqn45}
\end{equation}%
Define 
\begin{equation}
\theta _{n}=\frac{\delta ^{2\left( n+1\right) }\dprod\nolimits_{k=0}^{n}%
\left[ 1-a_{k}^{1}a_{k}^{2}\left( 1-\delta \right) \right] }{%
\dprod\nolimits_{k=0}^{n}\left[ 1-a_{k}^{0}\left( 1+\delta \right) \right] %
\left[ 1-a_{k}^{1}\left( 1+\delta \right) \right] \left[ 1-a_{k}^{2}\left(
1+\delta \right) \right] }\text{.}  \label{eqn46}
\end{equation}%
By the assumtion%
\begin{eqnarray}
&&\lim_{n\rightarrow \infty }\frac{\theta _{n+1}}{\theta _{n}}  \label{eqn47}
\\
&=&\lim_{n\rightarrow \infty }\frac{\frac{\delta ^{2\left( n+2\right)
}\dprod\nolimits_{k=0}^{n+1}\left[ 1-a_{k}^{1}a_{k}^{2}\left( 1-\delta
\right) \right] }{\dprod\nolimits_{k=0}^{n+1}\left[ 1-a_{k}^{0}\left(
1+\delta \right) \right] \left[ 1-a_{k}^{1}\left( 1+\delta \right) \right] %
\left[ 1-a_{k}^{2}\left( 1+\delta \right) \right] }}{\frac{\delta ^{2\left(
n+1\right) }\dprod\nolimits_{k=0}^{n}\left[ 1-a_{k}^{1}a_{k}^{2}\left(
1-\delta \right) \right] }{\dprod\nolimits_{k=0}^{n}\left[ 1-a_{k}^{0}\left(
1+\delta \right) \right] \left[ 1-a_{k}^{1}\left( 1+\delta \right) \right] %
\left[ 1-a_{k}^{2}\left( 1+\delta \right) \right] }}  \notag \\
&=&\lim_{n\rightarrow \infty }\frac{\delta ^{2}\left[
1-a_{n+1}^{1}a_{n+1}^{2}\left( 1-\delta \right) \right] }{\left[
1-a_{n+1}^{0}\left( 1+\delta \right) \right] \left[ 1-a_{n+1}^{1}\left(
1+\delta \right) \right] \left[ 1-a_{n+1}^{2}\left( 1+\delta \right) \right] 
}  \notag \\
&=&\delta ^{2}<1\text{.}  \notag
\end{eqnarray}%
It thus follows from ratio test that $\sum\limits_{n=0}^{\infty }\theta
_{n}<\infty $. Hence, we have $\lim_{n\rightarrow \infty }\theta _{n}=0$
which implies that the iterative sequence defined by (7) converges faster
than the iterative sequence defined by SP iteration method (6).

Using Noor iteration method (5), we get%
\begin{eqnarray}
\left\Vert \omega _{n+1}-u^{\ast }\right\Vert  &=&\left\Vert \left(
1-a_{n}^{0}\right) \omega _{n}+a_{n}^{0}T\varpi _{n}-u^{\ast }\right\Vert 
\label{eqn48} \\
&\geq &\left( 1-a_{n}^{0}\right) \left\Vert \omega _{n}-u^{\ast }\right\Vert
-a_{n}^{0}\left\Vert T\varpi _{n}-Tu^{\ast }\right\Vert   \notag \\
&\geq &\left( 1-a_{n}^{0}\right) \left\Vert \omega _{n}-u^{\ast }\right\Vert
-a_{n}^{0}\delta \left\Vert \varpi _{n}-u^{\ast }\right\Vert   \notag \\
&\geq &\left[ 1-a_{n}^{0}-a_{n}^{0}\delta \left( 1-a_{n}^{1}\right) \right]
\left\Vert \omega _{n}-u^{\ast }\right\Vert -a_{n}^{0}a_{n}^{1}\delta
^{2}\left\Vert \rho _{n}-u^{\ast }\right\Vert   \notag \\
&\geq &\left\{ 1-a_{n}^{0}-a_{n}^{0}\delta \left( 1-a_{n}^{1}\right)
-a_{n}^{0}a_{n}^{1}\delta ^{2}\left[ 1-a_{n}^{2}\left( 1-\delta \right) %
\right] \right\} \left\Vert \omega _{n}-u^{\ast }\right\Vert   \notag \\
&\geq &\left\{ 1-a_{n}^{0}-a_{n}^{0}\delta \left[ 1-a_{n}^{1}\left( 1-\delta
\right) \right] \right\} \left\Vert \omega _{n}-u^{\ast }\right\Vert   \notag
\\
&\geq &\left[ 1-a_{n}^{0}\left( 1+\delta \right) \right] \left\Vert \omega
_{n}-u^{\ast }\right\Vert   \notag \\
&\geq &\cdots   \notag \\
&\geq &\left\Vert \omega _{0}-u^{\ast }\right\Vert
^{n+1}\prod\limits_{k=0}^{n}\left[ 1-a_{k}^{0}\left( 1+\delta \right) \right]
\text{.}  \notag
\end{eqnarray}%
It follows by (42) and (47) that%
\begin{equation}
\frac{\left\Vert x_{n+1}-u^{\ast }\right\Vert }{\left\Vert \omega
_{n+1}-x_{\ast }\right\Vert }\leq \frac{\delta ^{2\left( n+1\right)
}\left\Vert x_{0}-u^{\ast }\right\Vert ^{n+1}\dprod\nolimits_{k=0}^{n}\left[
1-a_{k}^{1}a_{k}^{2}\left( 1-\delta \right) \right] }{\left\Vert \omega
_{0}-u^{\ast }\right\Vert ^{n+1}\prod\limits_{k=0}^{n}\left[
1-a_{k}^{0}\left( 1+\delta \right) \right] }\text{.}  \label{eqn49}
\end{equation}%
Define 
\begin{equation}
\theta _{n}=\frac{\delta ^{2\left( n+1\right) }\dprod\nolimits_{k=0}^{n}%
\left[ 1-a_{k}^{1}a_{k}^{2}\left( 1-\delta \right) \right] }{%
\prod\limits_{k=0}^{n}\left[ 1-a_{k}^{0}\left( 1+\delta \right) \right] }%
\text{.}  \label{eqn50}
\end{equation}%
By the assumption%
\begin{eqnarray}
\lim_{n\rightarrow \infty }\frac{\theta _{n+1}}{\theta _{n}}
&=&\lim_{n\rightarrow \infty }\frac{\delta ^{2}\left[
1-a_{n+1}^{1}a_{n+1}^{2}\left( 1-\delta \right) \right] }{\left[
1-a_{n+1}^{0}\left( 1+\delta \right) \right] }  \label{eqn51} \\
&=&\delta ^{2}<1\text{.}  \notag
\end{eqnarray}%
It thus follows from ratio test that $\sum\limits_{n=0}^{\infty }\theta
_{n}<\infty $. Hence, we have $\lim_{n\rightarrow \infty }\theta _{n}=0$
which implies that the iterative sequence defined by (7) converges faster
than the iterative sequence defined by Noor iteration method (5).
\end{proof}

By use of the following example due to \cite{QR}, it was shown in (\cite{CR}%
, Example 4.1) that CR iterative method \cite{CR} is faster than all of \
Picard (3), S \cite{S}, Noor (5) and SP (6) iterative methods for a
particular class of \ operators which is included in the class of
weak-contraction mappings satisfying (11). In the following, for the sake of
consistent comparison, we will use the same example as that of \ (\cite{CR},
Example 4.1) in order to compare the rates of convergence between the
Picard-S iterative scheme (7) and the CR iteration method \cite{CR} for
weak-contraction mappings. In the following example, for convenience, we use
the notations $\left( PS_{n}\right) $ and $\left( CR_{n}\right) $ for the
iterative sequences associated to Picard-S (7) and CR \cite{CR} iteration
methods, respectively.

\begin{example}
\cite{QR} Define a mapping $T:\left[ 0,1\right] \rightarrow \left[ 0,1\right]
$ as $Tx=\frac{x}{2}$. Let $a_{n}^{0}=a_{n}^{1}=a_{n}^{2}=0$, for $%
n=1,2,...,24$ and $a_{n}^{0}=a_{n}^{1}=a_{n}^{2}=\frac{4}{\sqrt{n}}$, for
all $n\geq 25$.

It can be seen easily that the mapping $T$ satisfies condition (11) with the
unique fixed point $0\in F_{T}$. Furthermore, it is easy to see that Example
1 satisfies all the conditions of Theorem 1. Indeed, let $x_{0}\neq 0$ be an
initial point for the iterative sequences $\left( PS_{n}\right) $ and $%
\left( CR_{n}\right) $. Utilizing Picard-S (7) and CR \cite{CR} iteration
methods we obtain 
\begin{eqnarray}
PS_{n} &=&\frac{1}{2}\left( \frac{1}{2}-\frac{4}{n}\right) x_{n}  \notag \\
&=&\cdots   \notag \\
&=&\dprod\limits_{k=25}^{n}\left( \frac{1}{4}-\frac{2}{k}\right) x_{0}\text{,%
}  \label{eqn52}
\end{eqnarray}%
\begin{eqnarray}
CR_{n} &=&\left( \frac{1}{2}-\frac{1}{\sqrt{n}}-\frac{4}{n}+\frac{8}{n\sqrt{n%
}}\right) x_{n}  \notag \\
&=&\cdots   \notag \\
&=&\dprod\limits_{k=25}^{n}\left( \frac{1}{2}-\frac{1}{\sqrt{k}}-\frac{4}{k}+%
\frac{8}{k\sqrt{k}}\right) x_{0}\text{.}  \label{eqn53}
\end{eqnarray}%
It follows from (51) and (52) that%
\begin{equation*}
\frac{\left\vert PS_{n}-0\right\vert }{\left\vert CR_{n}-0\right\vert }=%
\frac{\dprod\limits_{k=25}^{n}\left( \frac{1}{4}-\frac{2}{k}\right) x_{0}}{%
\dprod\limits_{k=25}^{n}\left( \frac{1}{2}-\frac{1}{\sqrt{k}}-\frac{4}{k}+%
\frac{8}{k\sqrt{k}}\right) x_{0}}
\end{equation*}%
\begin{equation*}
\text{ \ \ \ \ \ \ \ }=\dprod\limits_{k=25}^{n}\frac{\frac{1}{4}-\frac{2}{k}%
}{\frac{1}{2}-\frac{1}{\sqrt{k}}-\frac{4}{k}+\frac{8}{k\sqrt{k}}}
\end{equation*}%
\begin{equation*}
\text{ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }=\dprod\limits_{k=25}^{n}\frac{%
\left( k-8\right) \sqrt{k}}{2\left( k\sqrt{k}-2k-8\sqrt{k}+16\right) }
\end{equation*}%
\begin{eqnarray}
\text{ \ \ \ \ \ \ } &=&\dprod\limits_{k=25}^{n}\frac{\left( k-8\right) 
\sqrt{k}}{2\left( \sqrt{k}-2\right) \left( k-8\right) }  \notag \\
&=&\dprod\limits_{k=25}^{n}\frac{\sqrt{k}}{2\left( \sqrt{k}-2\right) }\text{.%
}  \label{eqn54}
\end{eqnarray}%
For all $k\geq 25$, we have%
\begin{eqnarray}
\frac{\left( k-2\right) \left( \sqrt{k}-4\right) }{4} &>&1  \notag \\
&\Rightarrow &\left( k-2\right) \left( \sqrt{k}-4\right) >4  \notag \\
&\Rightarrow &k\left( \sqrt{k}-4\right) >2\left( \sqrt{k}-2\right)   \notag
\\
&\Rightarrow &\frac{\sqrt{k}-4}{2\left( \sqrt{k}-2\right) }>\frac{1}{k} 
\notag \\
&\Rightarrow &\frac{\sqrt{k}}{2\left( \sqrt{k}-2\right) }<1-\frac{1}{k}\text{%
,}  \label{eqn55}
\end{eqnarray}%
which yields%
\begin{equation}
\frac{\left\vert PS_{n}-0\right\vert }{\left\vert CR_{n}-0\right\vert }%
=\dprod\limits_{k=25}^{n}\frac{\sqrt{k}}{2\left( \sqrt{k}-2\right) }%
<\dprod\limits_{k=25}^{n}\left( 1-\frac{1}{k}\right) =\frac{24}{n}\text{.}
\label{eqn56}
\end{equation}%
Therefore, we have%
\begin{equation}
\lim_{n\rightarrow \infty }\frac{\left\vert PS_{n}-0\right\vert }{\left\vert
CR_{n}-0\right\vert }=0\text{,}  \label{eqn57}
\end{equation}%
which implies that the Picard-S iterative scheme (7) is faster than the CR
iteration method \cite{CR}$.$
\end{example}

Having regard to R. Chugh et al.'s result (\cite{CR}, Example 4.1), L.B.
Ciric et al.'s results \cite{Ciric} and Example 1 above, we conclude that
Picard-S iteration method is faster than all Picard (3), Mann \cite{Mann},
Ishikawa \cite{Ishikawa}, S \cite{S}, Noor (5) and SP (6) iterative methods.

We are now able to establish the following data dependence result.

\begin{theorem}
Let $T$ with fixed point $u^{\ast }\in F_{T}\neq \emptyset $ be as in
Theorem 1 and $\widetilde{T}$ an approximate operator of $T$. Let $\left\{
x_{n}\right\} _{n=0}^{\infty }$ be an iterative sequence generated by (7)
for $T$ and define an iterative sequence $\left\{ \widetilde{x}_{n}\right\}
_{n=0}^{\infty }$ as follows%
\begin{equation}
\left\{ 
\begin{array}{c}
\widetilde{x}_{0}\in D\text{, \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \
\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ } \\ 
\widetilde{x}_{n+1}=\widetilde{T}\widetilde{y}_{n}\text{, \ \ \ \ \ \ \ \ \
\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ } \\ 
\widetilde{y}_{n}=\left( 1-a_{n}^{1}\right) \widetilde{T}\widetilde{x}%
_{n}+a_{n}^{1}\widetilde{T}\widetilde{z}_{n}\text{, \ \ \ \ \ \ } \\ 
\widetilde{z}_{n}=\left( 1-a_{n}^{2}\right) \widetilde{x}_{n}+a_{n}^{2}%
\widetilde{T}\widetilde{x}_{n}\text{, }n\in 
\mathbb{N}
\text{,}%
\end{array}%
\right.   \label{eqn58}
\end{equation}%
where $\left\{ a_{n}^{i}\right\} _{n=0}^{\infty }$, $i\in \left\{
1,2\right\} $ be real sequences in $\left[ 0,1\right] $ satisfying (i) $%
\frac{1}{2}\leq a_{n}^{1}a_{n}^{2}$ for all $n\in 
\mathbb{N}
$, and (ii) $\sum\limits_{n=0}^{\infty }a_{n}^{1}a_{n}^{2}=\infty $. If $%
\widetilde{T}\widetilde{u}^{\ast }=\widetilde{u}^{\ast }$ such that $%
\widetilde{x}_{n}\rightarrow \widetilde{u}^{\ast }$ as $n\rightarrow \infty $%
, then we have%
\begin{equation}
\left\Vert u^{\ast }-\widetilde{u}^{\ast }\right\Vert \leq \frac{%
5\varepsilon }{1-\delta }\text{,}  \label{eqn59}
\end{equation}%
where $\varepsilon >0$ is a fixed number.
\end{theorem}

\begin{proof}
It follows from (7), (11), (12), and (57) that%
\begin{eqnarray}
\left\Vert z_{n}-\widetilde{z}_{n}\right\Vert &\leq &\left(
1-a_{n}^{2}\right) \left\Vert x_{n}-\widetilde{x}_{n}\right\Vert
+a_{n}^{2}\left\Vert Tx_{n}-\widetilde{T}\widetilde{x}_{n}\right\Vert  \notag
\\
&\leq &\left[ 1-a_{n}^{2}+a_{n}^{2}\delta \right] \left\Vert x_{n}-%
\widetilde{x}_{n}\right\Vert +a_{n}^{2}L\left\Vert x_{n}-Tx_{n}\right\Vert
+a_{n}^{2}\varepsilon \text{,}  \label{eqn60}
\end{eqnarray}%
\begin{eqnarray}
\left\Vert y_{n}-\widetilde{y}_{n}\right\Vert &\leq &\left(
1-a_{n}^{1}\right) \delta \left\Vert x_{n}-\widetilde{x}_{n}\right\Vert
+a_{n}^{1}\delta \left\Vert z_{n}-\widetilde{z}_{n}\right\Vert  \notag \\
&&+\left( 1-a_{n}^{1}\right) L\left\Vert x_{n}-Tx_{n}\right\Vert
+a_{n}^{1}L\left\Vert z_{n}-Tz_{n}\right\Vert  \notag \\
&&+\left( 1-a_{n}^{1}\right) \varepsilon +a_{n}^{1}\varepsilon \text{,}
\label{eqn61}
\end{eqnarray}%
\begin{equation}
\left\Vert x_{n+1}-\widetilde{x}_{n+1}\right\Vert \leq \delta \left\Vert
y_{n}-\widetilde{y}_{n}\right\Vert +L\left\Vert y_{n}-Ty_{n}\right\Vert
+\varepsilon \text{.}  \label{eqn62}
\end{equation}%
From the relations (59), (60), and (61)%
\begin{eqnarray}
\left\Vert x_{n+1}-\widetilde{x}_{n+1}\right\Vert &\leq &\delta ^{2}\left[
1-a_{n}^{1}a_{n}^{2}\left( 1-\delta \right) \right] \left\Vert x_{n}-%
\widetilde{x}_{n}\right\Vert  \notag \\
&&+\left\{ a_{n}^{1}a_{n}^{2}\delta ^{2}L+\left( 1-a_{n}^{1}\right) \delta
L\right\} \left\Vert x_{n}-Tx_{n}\right\Vert  \notag \\
&&+L\left\Vert y_{n}-Ty_{n}\right\Vert +a_{n}^{1}\delta L\left\Vert
z_{n}-Tz_{n}\right\Vert  \notag \\
&&+a_{n}^{1}a_{n}^{2}\delta ^{2}\varepsilon +\left( 1-a_{n}^{1}\right)
\delta \varepsilon +a_{n}^{1}\delta \varepsilon +\varepsilon \text{.}
\label{eqn63}
\end{eqnarray}%
Since $a_{n}^{1}$, $a_{n}^{2}\in \left[ 0,1\right] $ $\ $and $\frac{1}{2}%
\leq a_{n}^{1}a_{n}^{2}$ for all $n\in 
\mathbb{N}
$%
\begin{equation}
1-a_{n}^{1}a_{n}^{2}\leq a_{n}^{1}a_{n}^{2}\text{,}  \label{eqn64}
\end{equation}%
\begin{equation}
1-a_{n}^{1}\leq 1-a_{n}^{1}a_{n}^{2}\leq a_{n}^{1}a_{n}^{2}\text{,}
\label{eqn65}
\end{equation}%
\begin{equation}
1\leq 2a_{n}^{1}a_{n}^{2}\text{.}  \label{eqn66}
\end{equation}%
Use of the facts $\delta $, $\delta ^{2}\in \left( 0,1\right) $, (63), (64),
and (65) in (62) yields 
\begin{eqnarray}
\left\Vert x_{n+1}-\widetilde{x}_{n+1}\right\Vert &\leq &\left[
1-a_{n}^{1}a_{n}^{2}\left( 1-\delta \right) \right] \left\Vert x_{n}-%
\widetilde{x}_{n}\right\Vert  \notag \\
&&+a_{n}^{1}a_{n}^{2}\left( 1-\delta \right) \left\{ \frac{L\delta \left(
1+\delta \right) \left\Vert x_{n}-Tx_{n}\right\Vert }{1-\delta }\right. 
\notag \\
&&\left. +\frac{2L\left\Vert y_{n}-Ty_{n}\right\Vert +2\delta L\left\Vert
z_{n}-Tz_{n}\right\Vert +5\varepsilon }{1-\delta }\right\} \text{.}
\label{eqn67}
\end{eqnarray}%
Define%
\begin{eqnarray}
\beta _{n} &:&=\left\Vert x_{n}-\widetilde{x}_{n}\right\Vert \text{,}
\label{eqn68} \\
\mu _{n} &:&=a_{n}^{1}a_{n}^{2}\left( 1-\delta \right) \in \left( 0,1\right) 
\text{,}  \notag \\
\gamma _{n} &:&=\frac{L\delta \left( 1+\delta \right) \left\Vert
x_{n}-Tx_{n}\right\Vert +2L\left\Vert y_{n}-Ty_{n}\right\Vert +2\delta
L\left\Vert z_{n}-Tz_{n}\right\Vert +5\varepsilon }{1-\delta }\geq 0\text{.}
\notag
\end{eqnarray}%
Hence, the inequality (66) perform all assumptions in Lemma 2 and thus an
application of Lemma 2 to (66) yields 
\begin{eqnarray}
0 &\leq &\lim \sup_{n\rightarrow \infty }\left\Vert x_{n}-\widetilde{x}%
_{n}\right\Vert  \label{eqn69} \\
&\leq &\lim \sup_{n\rightarrow \infty }\frac{L\delta \left( 1+\delta \right)
\left\Vert x_{n}-Tx_{n}\right\Vert +2L\left\Vert y_{n}-Ty_{n}\right\Vert
+2\delta L\left\Vert z_{n}-Tz_{n}\right\Vert +5\varepsilon }{1-\delta }. 
\notag
\end{eqnarray}%
We know from Theorem 1 that $\lim_{n\rightarrow \infty }x_{n}=u^{\ast }$ and
since $Tu^{\ast }=u^{\ast }$%
\begin{equation}
\lim_{n\rightarrow \infty }\left\Vert x_{n}-Tx_{n}\right\Vert
=\lim_{n\rightarrow \infty }\left\Vert y_{n}-Ty_{n}\right\Vert
=\lim_{n\rightarrow \infty }\left\Vert z_{n}-Tz_{n}\right\Vert =0\text{.}
\label{eqn70}
\end{equation}%
Therefore the inequality (68) becomes 
\begin{equation}
\left\Vert u^{\ast }-\widetilde{u}^{\ast }\right\Vert \leq \frac{%
5\varepsilon }{1-\delta }.  \label{eqn71}
\end{equation}
\end{proof}

\end{document}